Problem 4 (15 points) n January 1, 2012, Raiders Company had 100,000 shares of common stock outstanding an 20,000 shares of 8%, $100 par, cumulative preferred stock outstanding. The preferred stocks are convertible to 100,000 shares of common stocks. Raiders reported net income of $500,000 The income tax rate is 40%. Also outstanding at January 1, 2012 were fully vested incentive k options giving key employees the option to buy 40,000 common shares at $25. The arket price of the common shares averaged $40 during 2012 Raiders also had 4,000,706 rtible bonds outstanding throughout 2012. Each $1,000 bond is convertible into 100 of common stock. None of the bonds had been converted by December 31, 2012 ar stock options were exercised during the year Required: Compute basic and diluted earnings per share for Raiders Company for 2012.
Answer #1
Raiders Company
Basic EPS = (Net Income after taxes-preferred Dividend)/weighted no. of shares outstanding
= =(500000-2000000*8%)/100000
= 3.4
Diluted EPS = =620000/515000 (See working Notes below)
1.203883495
Net profit attributable attributable to common stockholders No. of equity shares Net profit attributable per shares
Net Income 340000 A 100000 3.4 Basic
Incentive stock Options 0 B 15000
Total 340000 C=A+B 115000 2.95652174 Dilutive
Convertible Bonds 280000 D 400000
Total 620000 E=C+D 515000 1.2038835 Dilutive
Preferred stock 160000 F 100000
Total 780000 G=E+F 615000 1.26829268 Anti-Dilutive
Working Note : Order of preference for dilution
Incentive stock option
Increase in earnings NIL
No. of incremental shares issued for no consideration 15000
(40000*(40-25)/40
EPS on increased share NIL/15000) NIL
Convertible Preference shares
Increase in earnings 160000
No. of incremental share 100000
EPS on increased share =160000/100000 1.6
Convertible Bonds
Increase in earnings = Interest 168000 =4000000*7%*(1-0.4)
No. of incremental share 400000
EPS on increased share =280000/400000 0.42
Note : as per above calculations, we can see that incentive options are most dilutive followed by convetible
bonds and convertible preference shares comes last in the order.
We appreciate rating of our answers. It really encourages us to improve or maintain quality of our answers.
Thank You